Ticket Number: HEA-NOC/20080109-7 Ticket Status: UPDATE
Ticket Type: unscheduled Resolver: HEA-NOC
Ticket Opened: 20080109 14:37 UTC Problem Start: 20080109 14:02 UTC
Ticket Update: 20080109 15:10 UTC Problem End: 20080109 14:58 UTC
Site/line: An Cheim
There is a fault with a switch in An Chéim which has caused the backup
paths between Institutes of Technology and An Chéim to fail. This also
revealed a config error which prevented the primary path operating
BGP sessions have been lost to the Eircom and Smart routers in An Cheim,
and CPE routers at Institutes of Technology have switched to Eircom/Smart
for paths to An Cheim.
Service between IoTs and An Cheim. The config error has been resolved and
the primary path is now being checked to make sure it's working.
20080109 14:31 DW Open ticket and investigating the circumstances of the
HRSP switchover. Also checking the status of the backup routes.
20080109 14:51 DW Fixed redistribution of routes, observed that HSRP
changed back to correct address on CPEs at 14:45, and alarms cleared a
short time later. Awaiting call back from HP to confirm whether
connectivity has been restored.
20080109 15:01 DW BGP sessions to Smart and Eircom routers restored at
20080109 15:07 DW Underlying cause was a problem with power supply to a
switch on site. Client reports service is restored. Ticket remains open to
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20080109-7
HEAnet Limited. Registered in Ireland, No. 275301.